# NeCo[Uncaptioned image]: Improving DINOv2’s spatial representations in 19 GPU hours with Patch Neighbor Consistency

### v1

@@ B.4 Extra Ablations
In 9(c), we vary the sorting steepness, denoted by $\beta$, for both teacher and student networks to evaluate the influence of hard or soft nearest neighbor assignments. The performance improves when the teacher’s steepness is higher or equal to the student’s, consistent with previous findings [13]. Our best results are achieved when both networks have equal steepness. However, extreme steepness values (e.g., 1000) harm performance. This is because sorting patch similarities lacks clear boundaries, and formulating it as a hard assignment can force incorrect orderings, negatively impacting performance.
Batch Size.
9(b) examines the impact of batch size on performance. We maintain the same experimental setup as reported in the paper for all ablations. As anticipated, smaller batch sizes result in marginal improvements due to the weaker supervisory signal. However, a moderately large batch size, such as the 64 used in the paper, shows significant improvements over the baseline. Furthermore, the trend suggests that increasing batch sizes could lead to even greater performance gains.
Number of Neighbors.
As mentioned in the method section, compute and sort the distances between each patch and all other patches using a differentiable sorting algorithm. Ultimately, The sorted distances for the same patches in different views should be similar. Here, we conduct an ablation study, selecting only the top $K$ distances for each patch instead of all distances. The results are reported in 9(a). As shown, incorporating more neighbours in NeCo  loss enhances performance. However, after reaching a threshold (e.g., 32 neighbors), adding more becomes less effective. This indicates our method’s robustness against this hyperparameter, not relying on computing distances for all batch patches to improve the baseline.
@@
We show the performance across different training epochs in 9(d). As the table shows, even after just one epoch of training, DINOv2R improves by 1% to 3% across various metrics. The performance continues to increase with more training epochs, but the improvements become smaller after 25 epochs, which is the number used in the paper.
Table 9: Ablations of the key parameters of our method. We evaluate the models by training a linear layer on top of the frozen representations (Lin.) or using the in-context (IC) evaluation of [10] using the validation images for PascalVOC12 and ADE20k.
(a) Num Neighbors
@@
Num Lin. IC Lin. IC
4 74.4 54.8 35.2 19.7
8 76.8 61.1 36.3 20.2
16 77.7 60.9 36.7 20.9
32 78.1 61.3 37.1 21.4
All 78.9 62.0 37.3 21.7
(b) Batch Size
@@
Batch Size Lin. IC Lin. IC
4 76.2 60.2 35.6 20.2
8 76.8 61.1 36.3 20.8
16 77.7 60.9 36.7 21.1
32 78.2 61.4 37.1 21.4
64 78.9 62.0 37.3 21.7
(c) Sorting steepness
@@
4 79.0 64.7 37.8 22.5 2h
8 80.7 65.9 39.6 23.2 5h
16 81.6 66.7 40.5 24.2 10h
25 81.9 66.9 40.8 24.4 19h
50 82.3 67.2 41.0 24.5 40h
B.5 Computational Analysis
We provide a detailed runtime analysis for DINO, CrIBo, TimeT, and NeCoin Table 10. All experiments are conducted on 8 NVIDIA RTX A6000-46GB GPUs. As shown in Table 3, starting from already existing pretrained networks, NeCo  only requires 2.5 GPU hours to enhance the performance of the state-of-the-art method, CrIBo, by 3.7% in linear segmentation. Furthermore, when NeCo  is applied to TimeT, starting from DINO initialization, it can surpass CrIBO with nearly 30% less total training time (20h + 1.6h + 2.5h vs. 70h). This trend is also observed in clustering and overclustering numbers, showing up to a 6% improvement over CrIBO when starting from TimeT, as detailed in Table 3. These results validate the computational efficiency of the proposed method.
Method Epoch time Init Num epoch Batch size Total training time Linear segmentation
@@ B.5 Computational Analysis
DINO 1:33 random 800 64 20h 43.9
TimeT 3:12 DINO 30 64 2h 58.2
CrIBo 5:16 random 800 64 70h 64.3
NeCo 4:48 TimeT 25 64 1.6h 64.8
NeCo 4:48 CrIBo 25 64 2.5h 68.0
@@ Appendix C Additional Visualizations
Figure 7: Fully unsupervised segmentation on Pascal for K=21. We extract foreground masks using the CBFE+CD method, following the approach outlined in [1]. These masks are then clustered into the number of objects present in the Pascal dataset, with $K=21$. As demonstrated, NeCo  yields distinct and accurate segmentation maps for each object.
@@ Appendix D Dataset Details
Pascal VOC 2012 [65] This dataset, the latest split version of trainaug, features 10,582 images and their annotations distributed across 21 classes, with one referring to the background class. The validation set consists of 1,449 images. Following [19] we ignore unlabelled objects as well as the boundary class. Moreover, for hyper-parameter tuning of the fully unsupervised segmentation method [1] that we apply on our method, we use the PVOC12 train split with 1464 images. Figure 3 shows the dataset images overlaid by the annotations.
@@
Pascal Context [52] This scene-centric dataset includes 4,998 training images covering 60 semantic classes, including the background. The validation set consists of 5,105 images. We use this dataset for the Linear Segmentation and Segmenter experiments, via the MMSegmentation Library [68].
@@
COCO-Stuff 164K [56] This scene-understanding dataset includes labels across 91 "stuff" categories and 80 "things" categories. The training set comprises 118,000 images, and the validation set contains 5,000 images. We follow the same setup as [1] and thus we use the COCO benchmark in two ways to isolate further the given object definitions.
@@
Concisely, we begin by extracting stuff annotations, which refer to objects without clear boundaries and often found in the background, using the COCO-Stuff annotations [56]. Then, we consolidate the 91 detailed labels into 15 broader labels, as described in [67] and we assign the general label “other” to non-stuff objects, as this label lacks specific semantic meaning. Non-Stuff objects are ignored during training and evaluation. We indicate this version of the dataset within our work as COCO-Stuff used in Overclusterring and Linear Segmentation in Appendix A.2.
@@
Next, we extract foreground annotations utilizing the panoptic labels from [70]. We combine the instance-level annotations into object categories using a script provided by the authors. Additionally, we consolidate the 80 detailed categories into 12 broad object classes.The background class is ignored during training and evaluation. This leads as to the COCO-Thing version of the dataset which we use for the Overclusterring and our Linear Segmentation in Appendix A.2.
@@
ADE20K [54] The dataset is a collection of images used for semantic segmentation tasks, featuring finely detailed labels across 150 unique semantic categories. Some of the categories include stuffs like sky and grass, as well as distinguishable objects like person, and a car. Overall, it includes a wide variety of scenes, with 20,210 images in the training set and 2,000 images in the validation set, making it one of the most challenging and diverse datasets for scene understanding. We use the full dataset in our experiments. In our experiments, we ignore the others label of the dataset.
@@
Imagenet [61] The dataset, is a large-scale visual database designed for use in visual object recognition research. It contains over 1.3 million images categorized into 1,000 object classes. Each image is labeled with detailed annotations, making it a critical resource for training and evaluating machine learning models, particularly in the field of computer vision. In our work, we also explore training on part of the Imagenet, the Imagenet100k that consists 100K images across 100 classes, from the original dataset.

### v2

@@ B.7 Extra Ablations
In 14(e), we vary the sorting steepness, denoted by $\beta$, for both teacher and student networks to evaluate the influence of hard or soft nearest neighbor assignments. The performance improves when the teacher’s steepness is higher or equal to the student’s, consistent with previous findings (Caron et al., 2021). Our best results are achieved when both networks have equal steepness. However, extreme steepness values (e.g., 1000) harm performance. This is because sorting patch similarities lacks clear boundaries, and formulating it as a hard assignment can force incorrect orderings, negatively impacting performance.
@@
We show the performance across different training epochs in 14(f). As the table shows, even after just one epoch of training, DINOv2R improves by 1% to 3% across various metrics. The performance continues to increase with more training epochs, but the improvements become smaller after 25 epochs, which is the number used in the paper.
Patch Similarity Metric.
We show the effect of using different metrics for computing pair-wise patch similarity in 14(c). As the table shows, Cosine similarity is consistently better than Euclidean distance. We have used cosine similarity in all our experiments.
ROI-Align Effect.
ROI-Align component can be removed if the multi-crop augmentation is omitted or by using the same Global Crop across the teacher and student branches. We ablate the effect of removing ROI align in 14(d). As shown by the results, this removal negatively affects performance due to using weaker self-supervised supervision through the augmentations.
Table 14: Ablations of the key parameters of our method. We evaluate the models by training a linear layer on top of the frozen representations (Lin.) or using the in-context (IC) evaluation of Balazevic et al. (2023) using the validation images for PascalVOC12 and ADE20k.
(a) Nearest-neighbour selection
@@
NN Lin. IC Lin. IC
intra 78.1 61.2 36.3 21.3
inter 78.9 62.0 37.3 21.7
(b) Sorting algorithm
@@
Method Lin. IC Lin. IC
 ✗ 47.3 17.8 15.8 5.1
Odd-even 78.9 62.2 37.0 21.6
Bitonic 78.9 62.0 37.3 21.7
(c) Patch similarity
Pascal ADE20K
Metric Lin. IC Lin. IC
Euc. 78.1 60.2 36.4 21.1
Cos. 78.9 62.0 37.3 21.7
(d) ROI Align
Pascal ADE20K
Lin. IC Lin. IC
 ✗ 75.8 60.2 35.6 19.3
 ✓ 78.9 62.0 37.3 21.7
(e) Sorting steepness
@@
4 79.0 64.7 37.8 22.4 2h
8 80.0 65.6 38.9 22.9 5h
16 80.6 66.1 39.5 23.3 10h
25 81.3 66.5 40.1 23.7 19h
50 81.6 66.7 40.2 23.8 40h
B.8 Computational Analysis
We provide a detailed runtime analysis for DINO, CrIBo, TimeT, and NeCo  in Table 15. All experiments are conducted on 8 NVIDIA RTX A6000-46GB GPUs. The results are reported on COCO-Things linear segmentation. The results show NeCo  significantly improves computational efficiency and performance. First, DINO and CrIBo are finetuned for 25 additional epochs starting from their existing checkpoints to match the extra training performed by NeCo . As the table shows, with the same number of extra epochs, NeCo  outperforms both models across all metrics, demonstrating that the improvement stems from the proposed loss function rather than extended training. Secondly, with only 2.5 GPU hours of extra training on top of CrIBo, NeCo  boosts its performance in linear segmentation by 3.7%. These results show that NeCo  not only enhances computational efficiency but also achieves superior results.
Method Dataset Epoch Time Init Epochs GPU Hours K=GT K=500 LS
@@ B.8 Computational Analysis
DINO ImageNet 15:33 Random 800 $\sim$8 days 5.4 19.2 43.9
DINO ImageNet 15:33 Random 800 + 25 $\sim$8 days + 6.5h 7.0 20.3 37.4
TimeT YTVOS 3:12 DINO 30 $\sim$8 days + 2h 18.4 44.6 58.2
NeCo COCO 4:48 DINO 25 $\sim$8 days + 1.6h 16.9 50.0 62.4
CrIBo ImageNet 20:37 Random 800 $\sim$11 days 14.5 48.3 64.3
CrIBo ImageNet 20:37 Random 800 + 25 $\sim$11 days + 9h 15.0 48.5 64.3
NeCo COCO 4:48 CrIBo 25 $\sim$11 days + 2.5h 21.1 54.0 68.0
@@ Appendix C Additional Visualizations
Figure 7: Fully unsupervised segmentation on Pascal for K=21. We extract foreground masks using the CBFE+CD method, following the approach outlined in Ziegler & Asano (2022). These masks are then clustered into the number of objects present in the Pascal dataset, with $K=21$. As demonstrated, NeCo  yields distinct and accurate segmentation maps for each object.
@@ Appendix D Dataset Details
Pascal VOC 2012 (Everingham et al., ) This dataset, the latest split version of trainaug, features 10,582 images and their annotations distributed across 21 classes, with one referring to the background class. The validation set consists of 1,449 images. Following Van Gansbeke et al. (2021) we ignore unlabelled objects as well as the boundary class. Moreover, for hyper-parameter tuning of the fully unsupervised segmentation method (Ziegler & Asano, 2022) that we apply on our method, we use the PVOC12 train split with 1464 images. Figure 3 shows the dataset images overlaid by the annotations.
@@
Pascal Context (Everingham et al., 2010) This scene-centric dataset includes 4,998 training images covering 60 semantic classes, including the background. The validation set consists of 5,105 images. We use this dataset for the Linear Segmentation and Segmenter experiments, via the MMSegmentation Library (MMSegmentation Contributors, 2020).
@@
COCO-Stuff 164K (Caesar et al., 2018) This scene-understanding dataset includes labels across 91 "stuff" categories and 80 "things" categories. The training set comprises 118,000 images, and the validation set contains 5,000 images. We follow the same setup as Ziegler & Asano (2022) and thus we use the COCO benchmark in two ways to isolate further the given object definitions.
@@
Concisely, we begin by extracting stuff annotations, which refer to objects without clear boundaries and often found in the background, using the COCO-Stuff annotations (Caesar et al., 2018). Then, we consolidate the 91 detailed labels into 15 broader labels, as described in Ji et al. (2019) and we assign the general label “other” to non-stuff objects, as this label lacks specific semantic meaning. Non-Stuff objects are ignored during training and evaluation. We indicate this version of the dataset within our work as COCO-Stuff used in Overclusterring and Linear Segmentation in Appendix A.2.
@@
Next, we extract foreground annotations utilizing the panoptic labels from Kirillov et al. (2019). We combine the instance-level annotations into object categories using a script provided by the authors. Additionally, we consolidate the 80 detailed categories into 12 broad object classes.The background class is ignored during training and evaluation. This leads as to the COCO-Thing version of the dataset which we use for the Overclusterring and our Linear Segmentation in Appendix A.2.
@@
ADE20K (Zhou et al., 2017) The dataset is a collection of images used for semantic segmentation tasks, featuring finely detailed labels across 150 unique semantic categories. Some of the categories include stuffs like sky and grass, as well as distinguishable objects like person, and a car. Overall, it includes a wide variety of scenes, with 20,210 images in the training set and 2,000 images in the validation set, making it one of the most challenging and diverse datasets for scene understanding. We use the full dataset in our experiments. In our experiments, we ignore the others label of the dataset.
@@
Imagenet (Russakovsky et al., 2015) The dataset, is a large-scale visual database designed for use in visual object recognition research. It contains over 1.3 million images categorized into 1,000 object classes. Each image is labeled with detailed annotations, making it a critical resource for training and evaluating machine learning models, particularly in the field of computer vision. In our work, we also explore training on part of the Imagenet, the Imagenet100k that consists 100K images across 100 classes, from the original dataset.
SPair-71k (Min et al., 2019) is a large-scale dataset of image pairs explicitly designed for semantic correspondence. It was constructed from images extracted from the well-known datasets of PASCAL (Everingham et al., 2010; Xiang et al., 2014). The task involves establishing correspondences at the pixel level between instances of objects in different images of the same class. There are 18 object categories, diversified between rigid and non-rigid objects, such as cars, aeroplanes, cats, and humans. Of these, 8 categories represent non-rigid objects, particularly challenging for semantic matching due to their deformability: cats, cows, humans, among others.
Each image is paired with its corresponding class-specific keypoints to help determine salient object parts. Furthermore, the pairs of images were annotated with a measure of viewpoint variation, describing how far the perspective can change between two images of the same category of an object. This annotation is done by human annotators for high-quality evaluation data. The combination of rich variability in object pose, appearance, and occlusion with background clutter, together with dense high-quality ground-truth correspondences, renders SPair-71k a uniquely challenging but comprehensive dataset to advance the frontiers of semantic correspondence and object matching.